# Function-Space MCMC for Bayesian Wide Neural Networks

### v1

@@ Abstract
Bayesian Neural Networks represent a fascinating confluence of deep learning and probabilistic reasoning, offering a compelling framework for understanding uncertainty in complex predictive models. In this paper, we investigate the use of the preconditioned Crank-Nicolson algorithm and its Langevin version to sample from the reparametrised posterior distribution of the weights as the widths of Bayesian Neural Networks grow larger. In addition to being robust in the infinite-dimensional setting, we prove that the acceptance probabilities of the proposed methods approach 1 as the width of the network increases, independently of any stepsize tuning. Moreover, we examine and compare how the mixing speeds of the underdamped Langevin Monte Carlo, the preconditioned Crank-Nicolson and the preconditioned Crank-Nicolson Langevin samplers are influenced by changes in the network width in some real-world cases. Our findings suggest that, in wide Bayesian Neural Networks configurations, the preconditioned Crank-Nicolson method allows for more efficient sampling of the reparametrised posterior distribution, as evidenced by a higher effective sample size and improved diagnostic results compared with the other analysed algorithms.
@@ 1 Introduction
Neural Networks (NNs) have become very popular in the recent years, due to the significant results achieved in various artificial intelligence tasks from speech recognition and image classification, to stock market prediction, healthcare and weather forecasting. Despite this widespread applicability and indisputable success, NNs are hindered by intrinsic shortcomings. The large number of parameters that makes NNs powerful function approximations, is also the reason why NNs are prone to overfitting and overconfidence in their predictions. Moreover, the black-box nature of NNs makes interpretability hard. Bayesian Neural Networks (BNNs) address some of these limitations by providing a principled way to incorporate uncertainty into complex predictive models. At its heart lies the framework of Bayesian Inference, specifically, the weights or parameters of the network are considered to be random variables, instead of deterministic quantities, whose prior beliefs are updated as data are gathered according to the Bayesian learning rule [1, 2, 3, 4, 5, 6, 7].
@@
Despite all these promising advantages and some recent progress, mainly under Gaussian initializations of the weights, BNNs have reached far less popularity than NNs, due to higher computational requirements and limited theoretical understanding. One of the major theoretical challenges concerns the comprehension of the parameter-space behavior of BNNs, in spite of the function one. We therefore explore sampling from the posterior distriution of wide BNNs, focusing on understanding its behavior and properties from a parameter-space perspective. In function space, in fact, under Gaussian initializations of the network’s weights, the distribution over the functions induced by wide BNNs has been proved to converge to the neural network Gaussian process (NNGP) limit [8, 9, 10, 11, 12, 13, 14]. Recently, [15] provided a counterpart of this result in the parameter space. Specifically, they proved that, after a reparametrization $\phi=T^{-1}(\theta)$ of the flattened NN weights $\theta\in\mathbb{R}^{d}$, the posterior distribution $p(\phi|\mathcal{D})$ converges in the KL-divergence to the standard Gaussian distribution $\mathcal{N}(0,I_{d})$ as the width of the BNN increases. This result not only gives a characterization of the wide-limit parameter distribution, but the closeness to the well-behaved standard normal distribution also suggests that an improvement in the mixing time of Markov chain Monte Carlo (MCMC) sampling procedures is possible. However, standard MCMC procedures are notably non-robust in the infinite-dimensional setting, since their proposal step needs to approach zero for the acceptance rate to remain constant, as the dimension of the space increases.
Figure 1: Comparison of the acceptance probability obtained using: i) the underdamped LMC algorithm (or Metropolis Adjusted Langevin Algorithm: MALA); ii) the pCN algorithm; iii) the pCNL method. The neural network architecture used is a fully-connected with one hidden layer, and the layer width varies among the following values: ${128,512,1024,2096,4192}$. The CIFAR-10 dataset is used, with the sample size fixed at $n=256$. The stepsize $\beta$ of the proposal step for all the methods is fixed at $0.1$. The acceptance rate of the pCN (orange line) increases steadily as the width of the BNN grows, suggesting improved performance in wide BNNs and highlighting its dimensional independence properties. In contrast, the LMC (blue line) initially shows a rapid acceleration in performance due to the effect of reparametrization. As the width of the network increases, the sampling distribution becomes more similar to an isotropic Gaussian distribution, which makes it better-behaved. However, as the width increases further, the sampler’s non-robustness in high-dimensional settings leads to a progressive deterioration in performance. This confirms the need to decrease the stepsize in order to maintain a constant acceptance rate, which introduces further correlations between the samples. The pCNL method (green line) shows a consistent acceptance rate across all dimensions, indicating a robustness to the dimensional scaling of the BNN. This stability may suggest that the pCNL algorithm effectively combines the strengths of the pCN and LMC methods.
@@
In this paper, we investigate and provide theoretical guarantees for the use of algorithms that are stable under mesh refinement, namely the preconditioned Crank-Nicolson (pCN) and Crank-Nicolson Langevin (pCNL), to sample from the reparametrised posterior distribution of the weights as the widths of BNNs grow larger. We show that the acceptance probability of both samplers, under reparametrization $T$, converges to 1 as the width of the BNN layers increases, independently of the chosen fixed stepsize. This result paves the way for a faster and more efficient sampling procedure in wide BNNs. Specifically, not only do the pCN and pCNL proposals avoid suffering from the curse of dimensionality, but the performance of the samplers intrinsically improves as the BNN becomes larger, without introducing any additional autocorrelation among the samplers. This lessens the need for longer burn-in periods and more samples for adequate convergence diagnostics, which affects standard MCMC methods.
@@
The theoretical guarantees for the use of the pCN and pCNL proposals in wide BNN configurations are empirically supported by a higher effective sample size and improved diagnostic results compared with the LMC algorithm. From a computational point of view, the pCN provides more scalability with respect to derivative-based MCMC algorithms and can scale well to large datasets and complex model. For this reason, the pCN method seems to effectively combine a small computational cost with a significant improvement in the quality of the collected samples, wherease the performancce of the pCNL procedure is not striking enough to justify its high computational cost. Notably, improvements in performance increase with the width size but are evident even far from the NNGP regime.
Figure 1 shows the potential of our approach. The acceptance probabilities of the underdamped Langevin Monte Carlo (LMC) [16] the pCN and the pCNL, applied to the posterior of the reparametrized weights, are plotted for a fully-connected network (FCN) with 1 hidden layer. In the samplers a fixed stepsize, $\beta=0.1$, in the proposal step has been used. As the width of the BNN grows large, the acceptance rate of the pCN algorithm steadily increases, whilst the one of LMC decreases as we go in higher dimension. The pCNL method, instead, shows a consistent acceptance probability across the dimensions. Worth-noting is the initial dramatic increase of the LMC acceptance rate, due to the reparametrization effect that bring the posterior close to the well-behaved standard Gaussian distribution. Further investigations will be provided in Section 4.1.
@@ 2.1 Notations and preliminaries
Let a dataset $\mathcal{D}={(\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n})}$, where $\mathbf{x}_{i}\in\mathbb{R}^{m}$ and $\mathbf{y}_{i}\in\mathbb{R}$. Let parameters $\theta$ belong to some parameter space $\Theta\subset\mathbb{R}^{D}$, for some $D\geq 1$. The prior distribution $p(\theta)$ represents our beliefs prior to observing the data, the beliefs are then updated accordingly to Bayes’ rule as data are gathered:
@@
where $p(\mathbf{y}|\theta,\mathbf{x})$ is the likelihood function.
@@
$g^{(l)}(\mathbf{x})=\psi(f^{(l)}(\mathbf{x})),\quad l=0,\dots,L,$ (1)
and
$f^{(l+1)}(\mathbf{x})=\frac{\sigma^{(l+1)}_{W}}{\sqrt{d^{(l)}}}g^{(l)}(\mathbf{x})W^{(l+1)}+\sigma^{(l+1)}_{b}\mathbf{1}^{l}b^{(l+1)},\quad l=0,\dots,L$
Thus, $\mathbf{y}=f^{L+1}(\mathbf{x})$ is the output of the network given the input $\mathbf{x}$. The (non-linear) activation function is denoted by $\psi$, $d^{(l)}$ denotes the width of the $l^{th}$ layer (with $d^{(0)}$ being the input size m) and $W^{(l)}\in\mathbb{R}^{d^{(l-1)}\times d^{(l)}}$ is its weight matrix. Without loss of generality, in the following we will not make explicit the presence of the bias $b^{(l)}\in\mathbb{R}^{d^{(l)}}$, though we include it in the weight matrix by adding a constant dimension to the outputs of the previous layer so to simplify the notation. The factor $\sigma^{l+1}_{W}/\sqrt{d^{l}}$ is to be attributed to the NTK parametrization [17], and it allows to consider an isotropic Gaussian prior for the weights, removing its usual dependence on the width. Precisely, we denote
@@
the flattened weights of the $l^{th}$ layer, bias included, assumed to be, a priori, independent standard Gaussian random variables. With $\theta\in\mathbb{R}^{D}$ we consider all the flattened weights $\theta=[\theta^{(l)}]_{l=0,\dots,L}\sim\mathcal{N}(0,I_{D})$. We assume a Gaussian likelihood $p(\mathbf{y}|\theta,\mathbf{x})\sim\mathcal{N}(f^{(L+1)}(\mathbf{x}),\sigma^{2}I_{D})$ for the observation variance $\sigma^{2}>0$. Moreover, we assume that $\left|\psi(x)\right|$ is bounded by a function of the form $\exp(C\|x\|^{2-\epsilon}-c)$, for some $C,c,\epsilon>0$; we refer to [12] for similar assumptions on the activation function.
We are interested in theoretical guarantees for sampling from the posterior distribution of the weights as the width grows. [15] proposed the following data-dependent reparametrization of the readout layer weights of BNNs:
@@
where, defining $\Psi=\frac{\sigma^{(L+1)}_{W}}{\sqrt{d^{(L)}}}g^{(L)}(\mathbf{x})$, we have
@@
The reparametrized posterior distribution, whose density function is denoted by $p(\phi|\mathcal{D})$, is shown to converge in the KL-divergence to a standard Gaussian $\mathcal{N}(0,I_{D})$ as the width goes to infinity. In particular, the convergence to a simple isotropic Gaussian suggests potential improvements in the mixing speed of MCMC procedures, compared to sampling from the notably arduous BNN posterior. Nevertheless, standard MCMC algorithms, such as the Random Walk Metropolis-Hastings Algorithm (RW-MH) and the Metropolis Adjusted Langevin Algorithm (LMC), are notoriously ill-suited for the infinite-dimensional setting and must be carefully re-tuned as the dimension increases to avoid degeneracy in the acceptance probability. Indeed, the optimal scaling properties of the RW-MH and LMC algorithms have been derived in [18] and [19], respectively. There, it is shown that the proposal movements need to vanish as the dimension of the parameter space grows unbounded for factorial target distributions. Consequently, the efficiency of the RW-MH and LMC algorithms vanishes as well.
@@ 2.2 Function-Space MCMC
Since our interest lies in analyzing the behavior of wide networks, we need methods that are robust in the infinite dimensional setting. A class of such MCMC algorithms have been early derived and discussed in [20]. In this section, we focus on the pCN and pCNL methods and briefly contrast them with their traditional counterparts: the RW-MH and the LMC algorithms. The slight modifications introduced in pCN and pCNL made them robust and well-defined in high dimensional settings, contrasting the well known degeneracy of the standard MCMC acceptance probability as the dimension increases.
@@
Instead, the pCN algorithm relies on the following modification of the (standard) random walk proposal:
@@
where $\mathcal{C}$ is the covariance operator of the prior Gaussian measure. See [20] and references therein for details on RH-MC and pCN proposals.
The pCN algorithm is reversible with respect to an arbitrary Gaussian measure, and hence admits such Gaussian measure as invariant distribution. Specifically, in our context, we set the invariant Gaussian measure to be equal to an isotropic Gaussian distribution (the prior). The acceptance probability reduces to:
@@
where $\ell$ is the log-likelihood. The steps of the pCN procedures are listed in Algorithm 1.
@@
3:     Propose $v=\sqrt{1-\beta^{2}}u^{(n)}+\beta N(0,C)$
@@
Despite being agnostic about which parts of the state space are more probable, the pCN proposal is, by construction, well-defined in the infinite-dimensional setting. This motivated us to study the procedure to sample from the above reparametrized weights posterior distribution. In our context, the proposal 4 is
$\phi^{*}=\sqrt{1-\beta^{2}}\phi+\beta w\quad w\sim\mathcal{N}(0,I_{D})\quad\implies\quad q(\phi^{*}|\phi)=\mathcal{N}(\sqrt{1-\beta^{2}}\phi,\beta^{2}I_{D}),$
@@
where $\beta\in[0,1)$. Now, by including this expression in the acceptance probability 5, we can prove the following theorem.
Theorem 2.1
@@
Consider the BNN model with the reparametrisation 2. Then, the acceptance probability of the pCN algorithm to sample from the reparametrised weight posterior, for any $\beta\in[0,1)$, converges to $1$ as the width of the network increases. If $d_{min}$ is the smallest among the network’s layer widths, then
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(\phi^{*}|\mathcal{D})q(\phi|\phi^{*})}{p(\phi|\mathcal{D})q(\phi^{*}|\phi)}\right\}\to 1\quad\text{as }\,d_{min}\to\infty$
See Appendix A for the proof of Theorem 2.1.
@@
The MALA algorithm incorporates gradient information from the log-posterior distribution in the proposal mechanism to guide the sampling process. More precisely, this method involves simulating Langevin diffusion such that the solution to the time evolution equation is a stationary distribution that equals the target density (in Bayesian contexts, the posterior distribution). This approach is particularly effective in exploring complex, high-dimensional probability distributions by making proposals that are more informed and therefore likely to be accepted. The mathematical expression of the proposal is:
@@
where $v$ is the current position of the algorithm in the parameter space, and $\epsilon$ is the step size, which is a tuning parameter that allows to control the scale of the updates. In particular, $\nabla\log p(\phi|D)$ is the gradient of the log-posterior density, which provides the direction towards higher probability densities.
While MALA provides an efficient way to explore the parameter space, it can still struggle in scenarios where the parameter space is high-dimensional. This led to the development of the pCNL algorithm, which introduces little modifications to better handle these challenges. Its proposal is derived by discretization of a stochastic partial differential equation (SPDE) which is invariant for target measure and is given by:
$v=\frac{1}{2+\delta}\left[(2-\delta)u+2\delta\mathcal{C}\mathcal{D}\ell(u)+\sqrt{8\delta}w\right],\quad w\sim\mathcal{N}(0,\mathcal{C}),\quad\delta\in(0,2)$ (6)
where $\mathcal{C}$ is the covariance operator of the Gaussian prior measure, $\ell$ is the log-likelihood, and $\mathcal{D}\ell$ denotes its Fréchet derivative. Note that the parameter $\delta$ is directly related to the pCN stepsize $\beta$ by the relationship $\beta^{2}=8\delta/(2+\delta)^{2}$. Throughout the paper, we refer to the stepsize as $\beta$, and we keep into account this relationship when dealing with the pCNL proposal. When the pCNL is used as a proposal for a MH scheme, the acceptance probability is given by
$a(v|u)=min\{1,exp(\rho(u,v))-\rho(v,u)\}$ (7)
, where
$\rho(u,v)=-\ell(u)-\frac{1}{2}\langle v-u,\mathcal{D}\ell(u)\rangle-\frac{\delta}{4}\langle u+v,\mathcal{D}\ell(u)\rangle+\frac{\delta}{4}\|\sqrt{\mathcal{C}}\mathcal{D}\ell(u)\|^{2}$
These modifications ensure that pCNL retains the efficiency of MALA in terms of informed proposals while enhancing its applicability and robustness in more challenging high-dimensional settings. The steps of the pCNL procedures are presented in Algorithm 2.

### v4

@@ Abstract
Bayesian Neural Networks represent a fascinating confluence of deep learning and probabilistic reasoning, offering a compelling framework for understanding uncertainty in complex predictive models. In this paper, we investigate the use of the preconditioned Crank-Nicolson algorithm and its Langevin version to sample from a reparametrised posterior distribution of the neural network’s weights, as the widths grow larger. In addition to being robust in the infinite-dimensional setting, we prove that the acceptance probabilities of the proposed algorithms approach 1 as the width of the network increases, independently of any stepsize tuning. Moreover, we examine and compare how the mixing speeds of the underdamped Langevin Monte Carlo, the preconditioned Crank-Nicolson and the preconditioned Crank-Nicolson Langevin samplers are influenced by changes in the network width in some real-world cases. Our findings suggest that, in wide Bayesian Neural Networks configurations, the preconditioned Crank-Nicolson algorithm allows for a scalable and more efficient sampling of the reparametrised posterior distribution, as also evidenced by a higher effective sample size and improved diagnostic results compared with the other analysed algorithms.
@@ 1 Introduction
Neural Networks (NNs) have become very popular in the recent years, due to the significant results achieved in various artificial intelligence tasks from speech recognition and image classification, to stock market prediction, healthcare and weather forecasting. Despite this widespread applicability and indisputable success, NNs are hindered by intrinsic shortcomings. The large number of parameters that makes NNs powerful function approximations, is also the reason why NNs are prone to overfitting and overconfidence in their predictions. Moreover, the black-box nature of NNs makes interpretability hard. Bayesian Neural Networks (BNNs) address some of these limitations by providing a principled way to incorporate uncertainty into complex predictive models. At its heart lies the framework of Bayesian Inference, specifically, the weights or parameters of the network are considered to be random variables, instead of deterministic quantities, whose prior beliefs are updated as data are gathered according to the Bayesian learning rule (Khan et al.,, 2018; Maddox et al.,, 2019; Osawa et al.,, 2019; Dusenberry et al.,, 2020; Daxberger et al.,, 2022; Izmailov et al.,, 2021; Neal,, 1996).
@@
Despite all these promising advantages and some recent progress, mainly under Gaussian initializations of the weights, BNNs have reached far less popularity than NNs, due to higher computational requirements and limited theoretical understanding. One of the major theoretical challenges concerns the comprehension of the parameter-space behavior of BNNs, in spite of the function one. We therefore explore sampling from the posterior distriution of wide BNNs, focusing on understanding its behavior and properties from a parameter-space perspective. In function space, in fact, under Gaussian initializations of the network’s weights, the distribution over the functions induced by wide BNNs has been proved to converge to the neural network Gaussian process (NNGP) limit (de G. Matthews et al.,, 2018; Lee et al.,, 2018; Novak et al.,, 2020; Garriga-Alonso et al.,, 2019; Yang,, 2021; Hron et al.,, 2020; Bracale et al.,, 2021). Recently, Hron et al., (2022) provided a counterpart of this result in the parameter space. Specifically, they proved that, after a reparametrization $\phi=T^{-1}(\theta)$ of the flattened NN weights $\theta\in\mathbb{R}^{d}$, the posterior distribution $p(\phi|\mathcal{D})$ converges in the KL-divergence to the standard Gaussian distribution $\mathcal{N}(0,I_{d})$ as the width of the BNN increases. This result not only gives a characterization of the wide-limit parameter distribution, but the closeness to the well-behaved standard normal distribution also suggests that an improvement in the mixing time of Markov chain Monte Carlo (MCMC) sampling procedures is possible. However, standard MCMC procedures are notably non-robust in the infinite-dimensional setting, since their proposal step needs to approach zero for the acceptance rate to remain constant, as the dimension of the space increases.
1.1 Our contributions
@@
In this paper, we investigate and provide some theoretical guarantees for the use of MCMC algorithms that are stable under mesh refinement, namely the preconditioned Crank-Nicolson (pCN) and Crank-Nicolson Langevin (pCNL), to sample from the reparametrised posterior distribution of the weights as the widths of BNNs grow larger. Our primary objective is to advance the theoretical understanding of inference in BNNs, rather than to introduce a new state-of-the-art algorithm for practical applications. We show that the acceptance probability of both samplers, under reparametrization $T$, converges to 1 as the width of the BNN layers increases, independently of the chosen fixed stepsize. This result paves the way for a faster and more efficient sampling procedure in wide BNNs. Specifically, not only do the pCN and pCNL proposals avoid suffering from the curse of dimensionality, but the performance of the samplers intrinsically improves as the BNN becomes larger, without introducing any additional autocorrelation among the samplers. This lessens the need for longer burn-in periods for adequate convergence diagnostics, which affects standard MCMC methods.
@@
The theoretical guarantees for the use of the pCN and pCNL proposals in the context of wide BNN configurations are empirically supported by a higher effective sample size and improved diagnostic results compared with the LMC algorithm. From a computational point of view, the pCN provides more scalability with respect to derivative-based MCMC algorithms and can scale well to large datasets and complex model. For this reason, the pCN method seems to effectively combine a small computational cost with a significant improvement in the quality of the collected samples, whereas the performance of the pCNL procedure is not striking enough to justify its high computational cost. Notably, the improvements in performance increase with the width size but are evident even far from the NNGP regime.
1.2 Organization of the paper
The paper is structured as follows. In Section 2 and Section 3 we present our methodologies based on the pCN and the pCNL proposals in wide BNNs. In Section 4 we provide an empirical validation of our methodologies. In Section 5 we discuss some directions for future research. The supplementary material contains the proofs of our results, and additional numerical experiments.
@@ 2.1 Notations and preliminaries
Let a dataset $\mathcal{D}={(\mathbf{x}_{1},\mathbf{y}_{1}),\dots,(\mathbf{x}_{n},\mathbf{y}_{n})}$, where $\mathbf{x}_{i}\in\mathbb{R}^{m}$ and $\mathbf{y}_{i}\in\mathbb{R}$. Let parameters $\theta$ belong to $\Theta\subset\mathbb{R}^{D}$, for some $D\geq 1$. The prior distribution $p(\theta)$ represents our beliefs prior to observing the data, the beliefs are then updated accordingly to Bayes’ rule as data are gathered:
@@
where $p(\mathbf{y}|\theta,\mathbf{x})$ is the likelihood function, i.e. the model for the data.
@@
and for $l=0,\dots,L$
$g^{(l)}(\mathbf{x})=\psi(f^{(l)}(\mathbf{x})),$
$f^{(l+1)}(\mathbf{x})=\frac{\sigma^{(l+1)}_{W}}{\sqrt{d^{(l)}}}g^{(l)}(\mathbf{x})W^{(l+1)}+\sigma^{(l+1)}_{b}\mathbf{1}^{l}b^{(l+1)}$
Thus, $\mathbf{y}=f^{L+1}(\mathbf{x})$ is the output of the network given the input $\mathbf{x}$. The (non-linear) activation function is denoted by $\psi$, $d^{(l)}$ denotes the width of the $l^{th}$ layer (with $d^{(0)}$ being the input size m) and $W^{(l)}\in\mathbb{R}^{d^{(l-1)}\times d^{(l)}}$ is its weight matrix. Without loss of generality, in the following we will not make explicit the presence of the bias $b^{(l)}\in\mathbb{R}^{d^{(l)}}$, though we include it in the weight matrix by adding a constant dimension to the outputs of the previous layer so to simplify the notation. The factor $\sigma^{l+1}_{W}/\sqrt{d^{l}}$ is to be attributed to the NTK parametrization (Jacot et al.,, 2020), and it allows to consider an isotropic Gaussian prior for the weights, removing its usual dependence on the width. Precisely, we denote
@@
the flattened weights of the $l^{th}$ layer, bias included, assumed to be, a priori, independent standard Gaussian random variables. With $\theta\in\mathbb{R}^{D}$ we consider all the flattened weights $\theta=[\theta^{(l)}]_{l=0,\dots,L}\sim\mathcal{N}(0,I_{D})$. We assume a Gaussian likelihood $p(\mathbf{y}|\theta,\mathbf{x})\sim\mathcal{N}(f^{(L+1)}(\mathbf{x}),\sigma^{2}I_{D})$ for the observation variance $\sigma^{2}>0$. Moreover, we assume that $\left|\psi(x)\right|$ is bounded by a function of the form $\exp(C\|x\|^{2-\epsilon}+c)$, for some $C,c,\epsilon>0$. This assumption still allows for functions that grow faster than exponential while remaining $L^{1}$ and $L^{2}$-integrable against Gaussian measures, with no constraint on their smoothness and encompasses almost all activations functions relevant in practice. See Yang, (2021) and references therein for similar assumptions on the activation function.
We are interested in theoretical guarantees for sampling from the posterior distribution of the weights as the width grows. Hron et al., (2022) proposed the following data-dependent reparametrization of the readout layer weights of BNNs:
@@
where, by defining the quantity $\Psi=\frac{\sigma^{(L+1)}_{W}}{\sqrt{d^{(L)}}}g^{(L)}(\mathbf{x})$, we have
@@
The reparametrized posterior distribution, whose density function is denoted by $p(\phi|\mathcal{D})$, is shown to converge in the KL-divergence to a standard Gaussian $\mathcal{N}(0,I_{D})$ as the width goes to infinity. In particular, the convergence to a simple isotropic Gaussian suggests potential improvements in the mixing speed of MCMC procedures, compared to sampling from the notably arduous BNN posterior. Nevertheless, standard MCMC algorithms, such as the Random Walk Metropolis-Hastings Algorithm (RW-MH) and the Metropolis Adjusted Langevin Algorithm (LMC), are notoriously ill-suited for the infinite-dimensional setting and must be carefully re-tuned as the dimension increases to avoid degeneracy in the acceptance probability. Indeed, the optimal scaling properties of the RW-MH and LMC algorithms have been derived in Roberts et al., (1997) and Roberts and Rosenthal, (1998), respectively. There, it is shown that the proposal movements need to vanish as the dimension of the parameter space grows unbounded for factorial target distributions. Consequently, the efficiency of the RW-MH and LMC algorithms vanishes as well.
@@ 2.2 Function-space MCMC
Since our interest lies in analyzing the behavior of wide networks, we need methods that are robust in the infinite dimensional setting. A class of such MCMC algorithms have been early derived in Cotter et al., (2013). In this section, we focus on the pCN and pCNL methods and briefly contrast them with their traditional counterparts: the RW-MH and the LMC algorithms. The slight modifications introduced in pCN and pCNL made them robust and well-defined in high dimensional settings, contrasting the well known degeneracy of the standard MCMC acceptance probability as the dimension increases.
@@
Instead, the pCN algorithm relies on the following modification of the (standard) random walk proposal, i.e.,
@@
where $\mathcal{C}$ is the covariance operator of the prior Gaussian measure. We refer to the work of Cotter et al., (2013), and references therein, for details on RH-MC and pCN proposals.
The pCN algorithm is reversible with respect to an arbitrary Gaussian measure, and hence admits such Gaussian measure as invariant distribution. Specifically, in our context, we set the invariant Gaussian measure to be equal to an isotropic Gaussian distribution (the prior). Accordingly, the acceptance probability reduces to:
@@
where $\ell$ is the log-likelihood. The pCN is in Algorithm 1.
@@
3:     Propose $v=\sqrt{1-\beta^{2}}u^{(n)}+\beta\eta,\quad\eta\sim N(0,C)$
@@
Despite being agnostic about which parts of the state space are more probable, the pCN proposal is, by construction, well-defined in the infinite-dimensional setting. This motivated us to study the procedure to sample from the above reparametrized weights posterior distribution. In our context, the proposal 3 is of the form
$\phi^{*}$ $=\sqrt{1-\beta^{2}}\phi+\beta w\quad w\sim\mathcal{N}(0,I_{D})$
$\implies$ $\quad q(\phi^{*}|\phi)=\mathcal{N}(\sqrt{1-\beta^{2}}\phi,\beta^{2}I_{D}),$
@@
where $\beta\in[0,1)$. Now, by including this expression in the acceptance probability 4, we can prove the following theorem.
Theorem 2.1.
@@
Consider the BNN model with the reparametrisation 1. Then, the acceptance probability of the pCN algorithm to sample from the reparametrised weight posterior, for any $\beta\in[0,1)$, converges to $1$ as the width of the network increases. If $d_{min}$ denotes the smallest among the network’s layer widths, then as $d_{min}\to\infty$
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(\phi^{*}|\mathcal{D})q(\phi|\phi^{*})}{p(\phi|\mathcal{D})q(\phi^{*}|\phi)}\right\}\to 1$
We provide a high-level sketch of the proof of Theorem 2.1. We refer to Appendix A.1 for the full proof of Theorem 2.1.
Sketch of Proof.
The acceptance probability for the Metropolis-Hastings step is given by:
$a(\phi^{*}|\phi)=\min\left\{1,\frac{p(\phi^{*}|\mathcal{D})q(\phi|\phi^{*})}{p(\phi|\mathcal{D})q(\phi^{*}|\phi)}\right\}.$
In particular, by making use of the reparametrization $\phi=T^{-1}(\theta)$, the posterior distribution $p(\phi|\mathcal{D})$ converges in KL-divergence to a standard Gaussian $\mathcal{N}(0,I_{d})$ in the infinite-width limit. Furthermore, the pCN proposal is designed to preserve Gaussian measures, meaning that:
$q(\phi^{*}|\phi)=\mathcal{N}\left(\sqrt{1-\beta^{2}}\phi,\beta^{2}I_{d}\right).$
Using this, the ratio of proposal densities simplifies to the following:
$\frac{q(\phi|\phi^{*})}{q(\phi^{*}|\phi)}=\exp\left(\frac{1}{2}(\|\phi^{*}\|^{2}-\|\phi\|^{2})\right).$
Along similar lines, we can express the posterior probability ratio in terms of the empirical NNGP kernel $\hat{K}_{\sigma^{2}}$, which converges to a constant in the infinite-width limit:
$\frac{p(\phi^{*}|\mathcal{D})}{p(\phi|\mathcal{D})}\propto\frac{\sqrt{\det(\hat{K}^{*}_{\sigma^{2}})}\exp\left(\frac{1}{2}y^{\top}(\hat{K}^{*}_{\sigma^{2}})^{-1}y\right)}{\sqrt{\det(\hat{K}_{\sigma^{2}})}\exp\left(\frac{1}{2}y^{\top}(\hat{K}_{\sigma^{2}})^{-1}y\right)}.$ (5)
Since $\hat{K}_{\sigma^{2}}\to K_{\sigma^{2}}$ in the limit, the ratio converges to 1, implying that:
$a(\phi^{*}|\phi)\to 1\quad\text{as}\quad d_{\min}\to\infty.$
We address here the choice of the white noise reference measure. A common assumption in infinite-dimensional sampling theory is that the reference Gaussian measure should have a trace-class covariance operator, as recently developed in the work of Sell and Singh, (2022), meaning that the sum of its eigenvalues is finite. However, in our work, we select a reference measure with an identity covariance operator $C=I$, leading to a white noise prior. From an intuitive standpoint, our approach does not rely on an explicit basis representation as in Cotter et al., (2013). Instead, we consider an alternative formulation where the posterior distribution itself converges to the base distribution in the limit of increasing width. To clarify this, let us consider a simple example.
Let $\mathcal{N}(n)$ denote a centered Gaussian distribution with identity covariance matrix in $n$ dimensions. When the proposal is constructed via pCN, it takes the following form:
$\theta^{\prime}\leftarrow\sqrt{1-\beta^{2}}\theta+\beta\epsilon,\quad\epsilon\sim\mathcal{N}(n),$ (6)
for $\beta\in(0,1]$. If $\theta\sim\mathcal{N}(n)$, then we also have $\theta^{\prime}\sim\mathcal{N}(n)$, ensuring that the proposal preserves the reference measure.
In contrast, when using the Metropolis Random Walk proposal:
$\theta^{\prime}\leftarrow\theta+\beta\epsilon,\quad\epsilon\sim\mathcal{N}(n),$ (7)
we observe that even if $\theta\sim\mathcal{N}(n)$, the new sample $\theta^{\prime}$ is not distributed as $\mathcal{N}(n)$. Accordingly, such a critical difference leads to distinct acceptance rate behaviors as $n\to\infty$.
@@
The LMC algorithm incorporates gradient information from the log-posterior distribution in the proposal mechanism to guide the sampling process. More precisely, this method involves simulating Langevin diffusion such that the solution to the time evolution equation is a stationary distribution that equals the target density (in Bayesian contexts, the posterior distribution). This approach is particularly effective in exploring complex, high-dimensional probability distributions by making proposals that are more informed and therefore likely to be accepted. The mathematical expression of the proposal is:
@@
where $v$ is the current position of the algorithm in the parameter space, and $\epsilon$ is the stepsize, which is a tuning parameter that allows to control the scale of the updates. In particular, $\nabla\log p(\phi|D)$ denotes the gradient of the log-posterior density function, which directly provides the direction towards higher probability densities.
While LMC provides an efficient way to explore the parameter space, it can still struggle in scenarios where the parameter space is high-dimensional. This led to the development of the pCNL algorithm, which introduces little modifications to better handle these challenges. More precisely, its proposal is derived by discretization of a stochastic partial differential equation (SPDE) which is invariant for target measure and is given by:
$v=\frac{1}{2+\delta}\left[(2-\delta)u+2\delta\mathcal{C}\mathcal{D}\ell(u)+\sqrt{8\delta}w\right],$ (8)
where $w\sim\mathcal{N}(0,\mathcal{C})$, $\delta\in(0,2)$, $\mathcal{C}$ is the covariance operator of the Gaussian prior measure, $\ell$ is the log-likelihood, and $\mathcal{D}\ell$ denotes its Fréchet derivative. We refer to Appendix C.1 for more details on Fréchet derivative. Note that the parameter $\delta$ is directly related to the pCN stepsize $\beta$ by the relationship $\beta^{2}=8\delta/(2+\delta)^{2}$. Throughout the paper, we refer to the stepsize as $\beta$, and we keep into account this relationship when dealing with the pCNL proposal. When the pCNL is used as a proposal for a MH scheme, the acceptance probability is given by
$a(v|u)=min\{1,\exp(\rho(u,v))-\rho(v,u)\},$ (9)
where
$\rho(u,v)$ $=-\ell(u)-\frac{1}{2}\langle v-u,\mathcal{D}\ell(u)\rangle$
$-\frac{\delta}{4}\langle u+v,\mathcal{D}\ell(u)\rangle+\frac{\delta}{4}\|\sqrt{\mathcal{C}}\mathcal{D}\ell(u)\|^{2}$
In particular, these modifications ensure that pCNL retains the efficiency of LMC in terms of informed proposals while enhancing its applicability and robustness in more challenging high-dimensional settings. The steps of the pCNL procedures are presented in Algorithm 2.